OpenStudy (anonymous):

Multiply seven-eighths times twenty-eight. thirty-two 4 and three-eighths one thirty-second twenty-four and one-half

OpenStudy (gahm8684):

72/32

OpenStudy (gahm8684):

or 2.40625

OpenStudy (gahm8684):

oooh im sorry it is 28 not 20/8

OpenStudy (haseeb96):

78^28

OpenStudy (gahm8684):

that would be 49/2

OpenStudy (gahm8684):

or 24.5

OpenStudy (haseeb96):

@ikram002P will give the answer because we are students and she is teacher of math :)

OpenStudy (anonymous):

It is 24 and 1/2

OpenStudy (ikram002p):

well its 7/8 * 28 could u calculate that ?

OpenStudy (ikram002p):

so 7/8 * 28 =196/8= 49/2 =24.5

OpenStudy (haseeb96):

teacher is teacher :)

OpenStudy (ikram002p):

:P

OpenStudy (haseeb96):

thumps up for yur work

OpenStudy (haseeb96):

:) we will always pray for u

OpenStudy (ikram002p):

^.^